OpenStudy (anonymous):

Simplify the following expression. 2x − 6y + 2z − 5x − 3y + 8z −3x − 9y + 10z −3x + 9y + 10z −3x + 9y − 10z 3x − 9y + 10z

OpenStudy (anonymous):

It equals a: -3x - 9y + 10z Explanation: You have: 2x - 6y + 2z - 5x - 3y + 8z What you have to do is simply. How? By combining like terms. So what terms do we have that we could combine? We have the 2x and the -5x You have the -6y and the -3y and you have the 2z + 8z Now combine them! 2x - 5x= -3x -6y - 3y= -9y 2z + 8z= 10z So you get: 2x - 6y + 2z - 5x - 3y + 8z= -3x - 9y + 10z